1.find four consecutive integers such that twice the smallest is 12 more than the largest
2.find three consecutive integers such that the sum of the first two integers is 24 more than the third integer.
3.find three consecutive odd integers such that the sum of the first and the second is 27 less than 3 times the third
4.find three consecutive even integers such that the sum of the smallest and twice the second is 20 more than the third

Answer by edjones(7311)

1)
x, x+1, x+2, x+3
2x=x+3+12
x=15
.
2)
x, x+1, x+2
x+x+1=x+2+24
2x+1=x+26
x=25
.
3)
x, x+2, x+4
x+x+2=3(x+4)-27
2x+2=3x+12-27
2x+2=3x-15
x=17
.
4)
x, x+2, x+4
x+2(x+2)=x+4+20
x+2x+4=x+24
3x+4=x+24
2x=20
x=10
.
Ed